OpenStudy (anonymous):

10.9% = 4.8(1-d)/27.5, solve for d

OpenStudy (lgbasallote):

.109 = 4.8 - 4.8d/27.5 2.9975 = 4.8 - 4.8d -1.8025 = -4.8d d = 0.37552083